OpenStudy (anonymous):

Solve the system. The system may have one solution no solutions, or an infinite number of solutions. Please show steps. 3x-y = -2 y = 3x-2

OpenStudy (diyadiya):

Substitute (3x-2) in place of y (since 3x-2=y) 3x-y=-2 3x-(3x-2)=-2 3x-3x+2=-2 +2=-2 ( this is not possible ) So there's no solution

OpenStudy (diyadiya):

@algebraprobs understood?

OpenStudy (anonymous):

I get it! wow, thank you soooo much!

OpenStudy (diyadiya):

Anytime :)